OpenStudy (anonymous):

A wheel totates at 100rpm A circular body is placed on the wheel concentrically to reduce the angular velocity to 50rpm in 2s the reduction in angular velocity takes place in n revelution find n?

OpenStudy (anonymous):

@CliffSedge @ajprincess @akash123 @Algebraic! @phi @ghazi

OpenStudy (anonymous):

Find 100rpm as revolutions per second.

OpenStudy (anonymous):

Same with the 5rpm to keep units consistent.

OpenStudy (anonymous):

I mean change rev/min to rev/sec.

OpenStudy (anonymous):

The slow down happens in 2s, and I'd rather not call that 1/30 of a minute...

OpenStudy (anonymous):

100rpm = 5/3 rps

OpenStudy (anonymous):

You can use the final and initial angular velocities and the 2s time interval to find angular acceleration.

OpenStudy (anonymous):

10 pie/6

OpenStudy (anonymous):

pie? Are you converting to radians?

OpenStudy (anonymous):

\[\omega = \frac{ 2\Pi }{ T }\]

OpenStudy (anonymous):

Ok, that's fine. You'll need to convert that back into number of revolutions at the end, remember. Anyway, when you get the angular acceleration, then you can use the first kinematics equation to get the angle it moves through. \[\large \theta =\omega_it+0.5\alpha t^2\]

OpenStudy (anonymous):

This is what I got for the angular acceleration. \[\large\alpha=\frac{\Delta\omega}{\Delta t} = \frac{\pi /6-10\pi /3}{2s} =-\frac{19}{12}s^{-2} \] Is that what you got too?

OpenStudy (anonymous):

\[\alpha = \frac{ 5\Pi }{ 6 }\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

\[\omega i = \frac{ 2 \Pi 100 }{ 60 }\] \[\omega f = \frac{ 2 \Pi 50 }{ 60 }\]

OpenStudy (anonymous):

Is the final angular velocity 5rpm or 50rpm?

OpenStudy (anonymous):

\[\theta = \frac{ 10 \Pi }{ 3 }\]

OpenStudy (anonymous):

Yup....i got the answer ...and i am Sorry it was a typo Final shuld be 50rpm

OpenStudy (anonymous):

Alright. And you can easily convert those radians back to number of revolutions, right? (also, I'm pretty sure this can be solved by keeping everything in units of revolutions all the way through; converting to radians shouldn't be necessary.)

OpenStudy (anonymous):

Thxx.....@CliffSedge

OpenStudy (anonymous):

Anytime.